The terminal side of an angle passes through the given ordered pairs below. For each angle, find the values of the six trigonometric ratios.
(2,-sqrt( 5 ))
sin⁡θ = ?
cosθ = ?
tanθ = ?
cscθ = ?
secθ = ?
cotθ = ?
Thanks!!

Answer by stanbon(75887)

The terminal side of an angle passes through the given ordered pairs below. For each angle, find the values of the six trigonometric ratios.
(2,-sqrt( 5 ))
Note:: That point is in QIV : x positive ; y negative
x = 2 ; y = -sqrt(5)
r = sqrt[2^2 + (sqrt(5))^2] = sqrt[4+5] = 3
-------
sin⁡θ = y/r = -sqrt(5)/3
cosθ = x/r = 2/3
tanθ = y/x = -sqrt(5)/2
----------------------
cscθ = r/y = 3/-sqrt(5) = (-3/5)sqrt(5)
---------------------
secθ = r/x = 3/2
-----------------
cotθ = x/y = 2/(-sqrt(5)) = -(2/5)sqrt(5)
------
Cheers,
Stan H.
---------------